# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ 3 Fundamentals of Agentic AI Systems
Table 2: Traditional AI Agents vs. Agentic AI Systems
@@
Autonomy Model × Reactive or deliberative (e.g., cognitive architectures) Goal-driven planning & adaptation
$\bullet$ Fixed action sequences or symbolic reasoning $\bullet$ Recursive self-improvement
Cognitive Foundation × Symbolic logic, FSMs, or cognitive models Foundation models (LLMs/LIMs)
$\bullet$ Hand-coded or structured KBs $\bullet$ Emergent reasoning
@@
Intelligence Scope × Narrow & task-specific Broad & compositional
$\bullet$ Single-domain focus $\bullet$ Cross-domain transfer
Reasoning Approach × Deterministic or multi-step symbolic Multi-step & contextual
$\bullet$ Rule-based or production rules $\bullet$ Chain-of-Thought (CoT)
@@
Collaboration × Isolated or limited coordination Role-specialized coordination
$\bullet$ Manual decomposition $\bullet$ Automated hierarchy
Temporal Context × Episodic, often stateless Persistent memory
$\bullet$ Session-bound or episodic memory $\bullet$ VectorDB/LTM
@@
Orchestration × Hard-coded or fixed workflows Dynamic meta-agents
$\bullet$ Sequential or predefined goals $\bullet$ Conflict resolution
Tool Utilization × Static or domain-specific tools Planned API invocation
$\bullet$ Handcrafted interfaces $\bullet$ Function calling
@@
Context Awareness × Bounded or heuristic context Memory-augmented
$\bullet$ Limited or cognitive context $\bullet$ RAG architectures
Learning Mechanism × Limited or rule-based learning Emergent learning
$\bullet$ Fixed rules or chunking $\bullet$ Fine-tuning/RL
Human-AI Interaction × Scripted or cognitive modeling Natural language interfaces
$\bullet$ Limited or human-like reasoning $\bullet$ Conversational adaptability
Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ SOAR  $\triangleright$ ACT-R $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
= Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
3.1 Traditional AI Agents vs Agentic AI
Traditional agents operate through predefined rules [46], heuristic workflows [47], or deterministic logic [48], and perform well in narrow, well-defined environments. Cognitive Agents [49], model human-like cognitive functions, such as integrating perception, memory, and decision-making, often using cognitive systems architectures like [50] or ACT-R [51]. According to a recent work [52], authors distinguish between 4 types of knowledge: Concepts, Skills, Processes, Motives for cognitive agents. A key difference between Agentic AI and such cognitive architecture approaches is that the latter were mostly applied to robot planning and control (so applies to ‘physical operational environments’) but Agentic AI applies more broadly, also to non-physical operational environments (which could be made to operate entities in physical also environments).
In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities. Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [30, 53]. (2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [54]. (3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [55].
This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [56]. For example, systems like AutoGen [57] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [58]. This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [53]. A comparison of traditional AI agents and Agentic AI systems is presented in Table 2. For the scope of this work, we rely on LLM-based Agentic AI systems.
@@ 3.2 Multi-Agentic AI System (AMAS) Architecture
AMAS represents an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [59]. These agents access a Shared Toolchain Interface [60] to execute code, perform searches, or interact with domain-specific APIs.
@@
Figure 3: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms). The modular design highlights adaptive intelligence and inter-agent synergy.
As shown in Figure 3), communication and coordination are facilitated through a Communication Middleware [61] that allows agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator [62] governs high-level planning and delegates subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory [14], which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface [63], enabling users to prompt, correct, or halt agent behavior.
To ensure accountability, a Trust and Audit module monitors agent actions, logs tool usage, and generates behavioral traces [64]. However, this modular and distributed structure introduces significant TRiSM [65] challenges. With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing [66].
Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information [67], especially when data traverses multiple agents or tools. Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration [68]. Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors. Below, we discuss the architecture of AMAS.
Language Model Core (Agent Brain). At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” [69]. The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access). It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats. In many agent frameworks, such as AutoGPT [70], Baby AGI [71] and GPT Engineer [72], the LLM governs the full control loop, orchestrating the overall system behavior.
Planning and Reasoning Module. To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals. This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting (ToT) [73], where the model performs intermediate reasoning before arriving at a final decision. Reasoning is also one aspect of cognitive functionalities [74]. Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making. Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations. Techniques like Reasoning and Acting (ReAct) [75] exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
Memory Module. Agentic AI systems integrate memory to maintain context across iterations. This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) [9]. Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search. By reintegrating past data into the LLM prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning. Effective memory management enables adaptive, learning-driven behavior.
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes. Acronyms: CoT = Chain-of-Thought reasoning, FSM = Finite State Machine, PDDL = Planning Domain Definition Language, DAG = Directed Acyclic Graph, DB = Database, REPL = Read–Eval–Print Loop, API = Application Programming Interface, KV = Key–Value store, CLI = Command Line Interface, SDK = Software Development Kit, MCP = Model Context Protocol.
@@
AutoGPT [76] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
@@
BabyAGI [71] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
@@
GPT Engineer [77] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
@@
LangGraph [78] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
@@
AutoGen [57] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
@@
MRKL [79] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
@@
Reflexion [80] GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
@@
MetaGPT [81] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
@@
Voyager [82] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
@@
WebVoyager [83] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
@@
HuggingGPT [84] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
@@
CAMEL [85] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
@@
ChatDev [86] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
@@
CrewAI [87] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
@@
AgentVerse [88] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
@@
OpenAgents [89] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI [90] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
@@
Semantic Kernel [91] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
@@
OpenAI Swarm [92] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
@@
OpenAI Agents SDK [93] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
@@
Strands Agents [94] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
@@
LlamaIndex Agents [95] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
Tool-Use Interface. To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface [96]. This layer allows the invocation of external tools such as web search, APIs, code interpreters, or databases. The available tools are specified in the agent prompt with command schemas. When the LLM determines a tool is needed, it emits a structured command, which is executed externally. The result is fed back into the LLM as a new observation. This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically. Tool-use frameworks like Modular Reasoning, Knowledge, and Language (MRKL) [79] illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router. Similarly, approaches like Toolformer [96] train the model to insert API calls in its text generation.
Perception and Environment Interface. For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation-action interface is essential [97]. Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process. Conversely, the agent’s chosen actions are executed within the environment, and the resulting state changes are returned to the agent as observations. This loop supports sense-plan-act cycles that continue until the task is completed or halted. In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
Integration and Autonomy. These modules together form a closed-loop architecture. The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results [29]. Each iteration updates the agent’s context, enabling self-refinement through dynamic feedback. This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
Table  LABEL:tab:agentic_architectures presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice.
4 Threats and Risks in LLM-based Agentic Multi-Agent Systems
In this section, we present the main threats in AMAS and also discuss a taxonomy of emergent risk classes.
Figure 4: Overview of Key Risks and Threats in Agentic AI Systems. The left panel outlines system-level threats unique to agentic architectures, including Autonomy Abuse, Persistent Memory, and Agent Orchestration. The right panel presents a taxonomy of emergent risk classes : Adversarial Attacks, Data Leakage, Agent Collusion, and Emergent Behaviors that arises from multi-agent interactions, memory persistence, and decentralized decision-making.

### v5

@@ 2 Fundamentals of Agentic AI Systems
Table 2: Comparison of Traditional AI Agents and Agentic AI Systems
@@
Autonomy Reactive with fixed action sequences Goal-driven with adaptive behavior
Cognitive Foundation Symbolic logic, finite state machines Foundation models (large language models)
@@
Domain Scope Narrow, task-specific applications Broad, cross-domain capabilities
Reasoning Mechanism Rule-based, deterministic inference Multi-step, contextual reasoning (chain-of-thought)
@@
Collaboration Isolated agents, manual task decomposition Role-specialized agents, automated coordination
Memory Architecture Episodic, session-bound storage Persistent memory (vector databases, long-term memory)
@@
Orchestration Hard-coded, predefined workflows Dynamic orchestration via meta-agents
Tool Integration Static, hand-crafted interfaces Dynamic planning with API/function calling
@@
Context Handling Bounded context with heuristic rules Memory-augmented retrieval (RAG)
Learning Paradigm Rule updates, fixed knowledge base Emergent learning (fine-tuning, reinforcement learning)
User Interaction Scripted, template-based dialogue Natural language understanding and generation
Representative Systems MYCIN [daniel1997cadiag], ELIZA [walters2025eliza], SOAR [laird1987soar], ACT-R [taatgen2006modeling] AutoGen [autogen_agents_tutorial], ChatDev [qian2023chatdev], MetaGPT [hong2023metagpt]
2.1 Traditional AI Agents vs. Agentic AI
Traditional agents rely on predefined rules, heuristics, or deterministic logic and typically perform well in narrow, well-specified environments [muscettola1998remote, hannebauer1999formal, govaerts2010development]. Cognitive agents extend this paradigm by modeling human-like functions (perception, memory, decision-making) through cognitive architectures such as SOAR and ACT-R [langley2006cognitive, laird1987soar, anderson1993act]; recent work further categorizes their knowledge into concepts, skills, processes, and motives [langley2025spatial]. However, these approaches were largely developed for physical operational settings (e.g., robotic planning and control), whereas Agentic AI generalizes autonomy to both physical and non-physical environments.
In contrast, Agentic AI systems leverage foundation models (primarily LLMs) to support adaptive, goal-directed behavior via (i) multi-agent coordination, (ii) persistent context through memory, and (iii) dynamic orchestration of roles and tool use [qian2023chatdev, hong2023metagpt, borghoff2025human, schneider2025generative]. This enables long-horizon task execution and cross-domain generalization beyond traditional reactive pipelines [de2025open]. Table 2 summarizes these differences; in this survey, we focus on LLM-based Agentic AI systems.
@@ 2.2 Multi-Agentic AI System (AMAS) Architecture
Figure 9: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms).
As shown in Fig. 9, AMAS coordinate multiple LLM-based agents through a Communication Middleware [weyns2009agent], a Task Manager/Orchestrator [langgraph2024], and a World Model/Shared Memory [guo2024largelanguagemodelbased] that stores system state and task artifacts. Human oversight is provided via a Human-in-the-Loop (HITL) interface [cranshaw2017calendar]. To support trustworthy operation, a Trust and Audit module records actions and tool usage [kaur2024building], while a Security Gateway and Privacy Management Layer enforce access control and reduce leakage across agents and tools [zhong2023towards, feretzakis2024privacy]. An Explainability Interface provides interpretable rationales for multi-agent decisions to support transparency and trust calibration [singh2025explainable].
Communication & Collaboration (Agent Communication and Coordination)
Beyond basic messaging, recent work formalizes AMAS communication through explicit protocols. Google’s Agent-to-Agent (A2A) [^7] A2A Protocol supports discovery, capability advertisement, delegation, and result exchange, improving interoperability and auditability. Related efforts such as the Agent Network Protocol (ANP) [^8] Agent Network Protocol take a network-centric view with standardized interfaces for coordination and trust establishment. From a TRiSM perspective, protocolized communication exposes control points for authentication, authorization, logging, and lifecycle governance. Emerging semantic-aware communication [guo2024survey, wang2025internet] can improve coordination by encoding intent, but introduces risks such as semantic spoofing and privacy leakage.
Cognitive/Reasoning Layer (LLM Core + Planning)
At the center of an AMAS, a foundation model (typically an LLM) acts as the decision controller, taking a user goal plus role and tool specifications as input and generating the next actions (e.g., messages, tool calls, or structured function outputs) [Saravia_LLM_Agents_2024]. In agent frameworks, this controller drives an iterative loop of planning, acting, and updating context based on tool results and feedback.
Planning and Reasoning Module
To achieve complex goals, AMAS employ planning mechanisms that decompose objectives into sub-tasks and interleave reasoning with execution and feedback. This can be implemented through prompting strategies (e.g., CoT/ToT) or through explicit plan–act–observe loops such as ReAct [yao2023reactsynergizingreasoningacting, Saravia_Prompt_Engineering_Guide_Web_2024]. The plan is typically revised as new observations arrive, enabling adaptive long-horizon behavior.
Memory Module
AMAS integrate memory to maintain context across steps and sessions. Short-term memory is maintained in the working context, while long-term memory stores salient facts and past events for later retrieval (often via embedding-based search) [yang2024multi]. Memory improves continuity and reduces repetition, but also increases exposure to stale context and poisoning risks.
Learning Module
Some AMAS include learning components that adapt behavior over time (e.g., from feedback, rewards, or task outcomes), such as updating policies, prompt templates, or routing strategies based on past performance. In multi-agent settings, learning can amplify errors through feedback loops and introduce risks such as drift, poisoning, and emergent misalignment, motivating continuous monitoring and gated updates under TRiSM controls.
Data Storage & Retrieval
AMAS typically combine structured stores for system state and task metadata with vector-based retrieval to support RAG and long-term memory. Because data may be shared across agents and tools, storage and retrieval require strong access control, encryption, and audit logging to reduce leakage and support accountability [he2024security].
Tool-Use Interface
To extend capabilities beyond text generation, AMAS agents use a tool-use interface that enables structured calls to external systems (e.g., web search, APIs, code execution, and databases) [schick2023toolformerlanguagemodelsteach]. Tools and call schemas are specified in the agent configuration; when needed, the LLM emits a structured command that is executed externally, and the returned output is incorporated as a new observation for subsequent reasoning. Tool-routing frameworks such as MRKL [karpas2022mrklsystemsmodularneurosymbolic] exemplify this design by dispatching requests to specialized expert modules under an LLM controller, while training-based approaches (e.g., Toolformer) teach models to generate API calls during generation [schick2023toolformerlanguagemodelsteach].
Perception and Environment Interface
For agents interacting with dynamic environments (e.g., web interfaces, simulated worlds, or physical systems), an observation–action interface is essential [liu2025agent]. Perception modules translate raw inputs (sensor streams, images, or textual state) into representations the LLM can process; the agent’s selected actions are then executed in the environment, and resulting state changes are returned as new observations. This loop supports iterative sense–plan–act behavior until a task is completed or halted. In multimodal or robotic settings, additional perception models (e.g., vision encoders) may be included, while the core interaction pattern remains the same.
Monitoring and Governance Layer
A monitoring and governance layer provides system-wide oversight by collecting logs, traces, and metrics, enforcing policy constraints, and triggering alerts or interventions when anomalous behavior is detected. From a TRiSM perspective, this layer enables auditability and lifecycle governance across agent interactions, tool calls, and memory updates.
Integration and Autonomy
Together, these modules form an integrated architecture in which the LLM coordinates planning, tool use, and memory to act in an environment and update context from feedback [acharya2025agentic]. Table LABEL:tab:agentic_architectures summarizes representative systems and maps them to five design axes (LLM core, planning/reasoning, memory, tool use, and environment interface).
Key Insights
Overall, AMAS combine planning, tool use, memory, and environment interaction into a closed-loop system, which expands the attack surface and complicates accountability. Protocolized communication and monitoring create practical control points for authentication, logging, and governance, motivating TRiSM-oriented analysis for multi-agent deployments.
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
@@
AutoGPT [yang2023auto] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
@@
BabyAGI [nakajima_babyagi_archive_2024] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
@@
GPT Engineer [gpt-engineer] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
@@
LangGraph [wang2024agent] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
@@
AutoGen [wu2023autogen] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
@@
MRKL [karpas2022mrklsystemsmodularneurosymbolic] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
@@
Reflexion [shinn2023reflexionlanguageagentsverbal] GPT-3.5/4 Retry-reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
@@
MetaGPT [hong2024metagpt] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
@@
Voyager [wang2023voyageropenendedembodiedagent] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
@@
WebVoyager [he2024webvoyager] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
@@
HuggingGPT [shen2023hugginggpt] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
@@
CAMEL [li2023camel] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
@@
ChatDev [qian2024chatdevcommunicativeagentssoftware] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
@@
CrewAI [crewAI] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
@@
AgentVerse [chen2023agentverse] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
@@
OpenAgents [xie2023openagentsopenplatformlanguage] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI [superagi] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
@@
Semantic Kernel [microsoft_semantickernel] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
@@
OpenAI Swarm [openai_swarm] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
@@
OpenAI Agents SDK [openai_agents_sdk] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
@@
Strands Agents [strands_agents] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
@@
LlamaIndex Agents [llamaindex_agents] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
3 Threats and Risks in Agentic AI
This section summarizes major threats in Agentic AI and organizes emergent AMAS risks into a compact taxonomy.